OpenStudy (anonymous):

Someone help to find sin theta, cos theta, and tan theta for (11pi)/2

OpenStudy (anonymous):

yeah do you know how to find \(\frac{11\pi}{2}\) on the unit circle?

OpenStudy (anonymous):

I have the unit circle, but \[\frac{ 11\Pi }{ 2 }\] is not on there

OpenStudy (anonymous):



OpenStudy (anonymous):

keep counting out \(\frac{\pi}{2}\) until you get to 11 of them

OpenStudy (anonymous):

ok, so then you take the first of the ordered pair for sin

OpenStudy (anonymous):



OpenStudy (anonymous):

you will see that \(\frac{11\pi}{2}\) is coterminal with \(\frac{3\pi}{2}\)

OpenStudy (anonymous):

yes i found that, but then how do you find sin theta

OpenStudy (anonymous):

ordered pair at \(\frac{3\pi}{2}\) is \((0,-1)\) read off your answers from there

OpenStudy (anonymous):

so is the answer just 0?

OpenStudy (anonymous):

first coordinate is cosine second coordinate is sine

OpenStudy (anonymous):

ok so it is -1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks! you just saved me so much time on my homework.